OpenStudy (wilder.monday):

simplify 4x^2 + 12x - 16/2x + 10 all over 6x + 24/x^2 + 9x + 20 @ samanthagreer

OpenStudy (samanthagreer):

two things fail tag and close the question

OpenStudy (wilder.monday):

I closed it

OpenStudy (wilder.monday):

I started to simplify the denominators but I don't think I'm right

OpenStudy (samanthagreer):

first simplify the top part 4x^2+12x-16/2x+10 factor 4x^2+12x-16= (4x-16)(x-1) then do the bottom 2x+10= 2(x+5) then you get (4x-16)(x-1)/2(x+5) for the first part then simplify the second part 6x+24/x^2+9x+20 Simplify 6x+24= 6(x+4) factor x^2+9x+20= (x+5)(x+4) then you get 6(x+4)/(x+5)(x+4) for the second part then put it all together and simplify: (4x-16)(x-1) 6(x+4) ------------ /---------- 2(x+5) (x+5)(x+4) cancel out the (x+4) bc there are two of them and one is on the top and one is on the bottom then flip over the second fraction bc you are dividing (4x-16)(x-1) (x+5) ------------ /---------- 2(x+5) 6 then cancel out the (x+5) bc there are 2 of them and one is on the top and one is on the bottom then your final answer is (4x-16)(x-1)/12 because you are multiplying all of your top terms together and all of your bottoms terms together :)

OpenStudy (wilder.monday):

those aren't any of my choices though so...

OpenStudy (samanthagreer):

sorry

OpenStudy (wilder.monday):

its ok how about the last one?

OpenStudy (samanthagreer):

what about the last one

OpenStudy (wilder.monday):

can you help me with it?

OpenStudy (samanthagreer):

ooohh ok yeah just post then close